OpenStudy (anonymous):

4x^3-32 need to factor

OpenStudy (anonymous):

take 4 common and solve

OpenStudy (anonymous):

so it becomes 4 (x^3-8)

jhonyy9 (jhonyy9):

4(x^3 -2^3)

OpenStudy (anonymous):

now 8 can be written as 2^3 ....just simply and use a^3-b^3

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4(x^2 - 8) }\) can be factored(difference of cubes)

Parth (parthkohli):

4(x^3 - 8)*

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^3 - b^ 3 = (a - b)(a^2 + ab + b^2) }\)

OpenStudy (anonymous):

the both of you guys really confused me here

OpenStudy (btaylor):

1. Take the gcf of 4 out. \[\rightarrow 4(x^3 - 8)\] 2. Since 8 = 2^2, use the difference of squares (@ParthKohli gives the form correctly) to factor it.\[\rightarrow 4(x-2)(x^2 +2x+4)\]That is as factored as you can make it.

OpenStudy (anonymous):

thanks